OpenStudy (anonymous):

Which of the following pairs of coordinates represents the same point in polar coordinates? Check all that apply. A. (r, theta) and (r, theta + pi) B. (r, theta) and (-r, theta + 2pi) C. (r, theta) and (r, theta + 2pi) D. (r, theta) and (-r, theta)

OpenStudy (anonymous):

I am not too familiar with these but I think, if I remember correctly, that the answers would be A and C?

OpenStudy (anonymous):

Yep A and C *gives medal*

AZ:

To find the equivalent polar coordinates \((r, \theta) = (r, ~ \theta \pm 2n\pi)\) or \((r, \theta) = (-r, ~\theta \pm (2n+1)\pi)\) A simpler way to see this would be when n = 1 \((r, \theta) = (r, ~ \theta + 2\pi)\) \((r, \theta) = (-r, ~\theta + \pi)\) Trying to explain how this works would result in me writing a massive wall of text which would just be difficult to understand. This short video should help you visualize and understand it clearly. https://www.youtube.com/watch?v=-P_2tXNqwZQ

AZ:

One last thing I would like to add is that answer choice A is only correct if the "-" was omitted accidentally as it should read (-r, theta + pi)